# RAGServe: Fast Quality-Aware RAG Systems with Configuration Adaptation

### v1

@@ 4 RAGServe: Enabling per-query configuration adaptation for RAG
    Pruning configuration space: For each query, it first reduces the RAG configuration space to a smaller yet promising one that still yields high generation quality based on the profile of the query.
@@
    RAG scheduler: It then chooses the best configuration jointly with available system resource to reduce the delay while preserving quality.
The overall pipeline for RAGServe (Figure 6) is as follows: on the arrival of a RAG query, we prune the configuration space by first estimating the query’s profile (§4.1) and then mapping the profile to a pruned configuration space (§4.2). This reduces the possible configurations from 1500 down to about 50. With the pruned configuration space for the query, RAGServe’s scheduler then choose the best configuration for the query that achieves the best quality-latency trade-off based on the available system resource (§4.3).
Once the configuration is chosen, the RAG system executes the query using the chosen configuration—retrieving the selected number of chunks and uses the selected synthesis method to feed the retrieved chunks in the LLM’s input. RAGServe focuses on achieving the best quality-latency trade-off by choosing an appropriate number of chunks to retrieve.
Figure 6: RAGServe consists of a RAG controller which performs configuration space pruning and joint scheduling.
Figure 7: RAGServe RAG configuration selection workflow.
4.1 Estimating a query’s profile
Query profile:  To choose the correct RAG configurations, the first step of RAGServe is to create the profile of the query (as we see in Figure 7) by querying an LLM (we call this LLM query profiler). We ask the query profiler to estimate four high-level dimensions for each query.
@@ 4.1 Estimating a query’s profile
    Pieces of information required refers to the distinct, standalone pieces of information required to fully answer the query (e.g., the annual income from how many years is required to draw the trend of annual income). The output for this dimension is a number from 1-6.
@@
We should note that RAGServe is not the first to use query profile as a metric for deciding RAG configurations. Methods like AdaptiveRAG [31] have used LLM’s to estimate query profile but they only focus on one dimension (the number of chunks to retrieve).
Why the query profile could be estimated:  Estimating the aforementioned query profile is feasible, not only because of the reasoning power of LLMs [^2] We have tested both GPT-4o and Llama-70b as the profile query-profiler, and they yield similarly impressive results (§7). in analyzing natural language queries, but also because we provide sufficient information to the LLM-based profiler. RAGServe feeds the profile estimator with not only the query, but also a metadata of the database that contains the background document.
The metadata is a short description about the type of content in the database and its data size (chunk_size). Specifically, we use a single-line summaries already attached to the original source datasets as the metadata of the dataset. For example, the metadata for the KG RAG Finsec’s database  [45] contains quarterly financial reports and questions of Fortune 500 companies with a chunk_size of 1000. It describes the content topics of the chunks with information such as revenue growth indicators, product release information, sales etc.,. When presented with a query on financials of such a company, the LLM can use the metadata to decide questions like how much to summarize and how much reasoning is required. We give the details of the used metadata and how we generate the metadata in §A.
@@
Moreover, we observed that extra information does not significantly improve the profiler’s estimates. For instance, in theory, it helps to know the embedding algorithm and the LLM used by RAG. Yet, the embedding model and the LLM perform similarly overall across queries and datasets under our consideration. This explains their limited contribution to the profiler, though more future work will be needed to understand the wide implications.
4.2 Mapping query profile to reduced RAG configuration space
First, RAGServe obtains the query profile using the LLM. It then performs rule-based mapping to generate values for RAG configuration knobs (synthesis_method, num_chunks, and intermediate_length introduced in §2). To do so, we perform a subsequent rule-based mapping from the query profiler’s outputs to RAG configuration knobs.
How the profile helps:  To understand the importance of the query profiles further, let’s consider the following examples:
  • $\bullet$
    “Who is the current CEO of NVIDIA?” This query is not complex and does not require joint reasoning. Due to the query being simple with no reasoning required and one piece of information (name of CEO), this query can be mapped to map_rerank as we should only need a few chunks (1-2).
  • $\bullet$
    “Which month had the highest NVIDIA’s stock price the six months from January to June 2024?” This query is simple but needs to read information jointly (different stock prices). We need six pieces of information (stock price for every month). Due to the need for multiple pieces with reasoning, this query can be mapped to stuff with at least 6 chunks.
  • $\bullet$
    “What might be the reasons for NVIDIA’s month-on-month stock price change from January to June 2024”
    This query is complex, needs a lot of information, and needs to read information jointly. We need multiple pieces of information (stock prices, reasons for change, revenue, etc.) As multiple reasons need to be analyzed and compared here, summarizing all of the information first helps narrow it down to relevant information.
    Due to the needing multiple pieces with deep reasoning on why the prices changed, this query can be mapped to map_reduce with summarization with at least 6 chunks.
@@ 4.2 Mapping query profile to reduced RAG configuration space
10num_chunks = [Pieces of information , 3$\times$ Pieces of information]
@@
Algorithm 1 outlines this rule-based mapping process. This mapping is significantly helpful, it improves upon raw profiler outputs and converts them to usable RAG configurations. We use query complexity and whether joint reasoning is needed to determine the synthesis_method. Simple queries that don’t need any reasoning can answered with map_rerank while queries that require joint reasoning need stuff or map_reduce.
We decide the range of synthesis_method selections based on two of the profile dimensions estimated in §4.1, i.e., the “Query complexity” and the “Joint reasoning requirement”. We then decide the range of values for num_chunks based on the profile dimension of the “Pieces of information required”, i.e., $n$—specifically, we set the range of num_chunks to be $1-3$ times of $n$. We do not directly set num_chunks at $n$, because it (1) gives some leeway for the retrieval logic (e.g., typically Bert-embedding-based) [^4] A typical RAG retriever these days will have to retrieve 2-3$\times$ more chunks than minimally required to provide sufficient information for the LLM inference [24, 49]. to find necessary information, and (2) provides the room for the scheduler to select the configuration that fits in available memory. Finally, we get the intermediate_length range directly from the “summary length” estimate, which is already a value range.
Figure 8: RAGServe joint schedules RAG configurations with available GPU memory
4.3 Joint configuration-scheduling adaptation
@@
Once provided with the narrowed range of each RAG configuration knob (synthesis_method, num_chunks and intermediate_length), we need to choose the actual RAG configuration, which is aware of the current system resource (GPU memory). If we pick configurations which do not fit in current memory it will lead to additional delay, by waiting for the GPU memory to free up.
@@
We have RAGServe’s pruned configuration space where the quality is high, we now focus on choosing the best configuration which fits in memory, without focusing on quality.
@@
Why do we need to choose the scheduling jointly?  We motivate the need for joint scheduling along with the RAG configuration choice in Figure 8.
@@
Consider a setup where we tune only one RAG configuration knob of synthesis_method. Other knobs num_chunks and intermediate_length are fixed at 20 and 100 respectively. Let’s assume both stuff and map_reduce are present in the pruned space. For the scheduling knob, we consider the amount of GPU memory available in the current batch.
@@
The baseline system in Figure 8 (a) doesn’t consider other jobs in the system and does not evaluate the amount of available resource to make its scheduling decision. Due to its long input length with 20 chunks, stuff turns out to be memory-intensive. If the available GPU memory is low, stuff doesn’t fit in memory and needs to be queued. This ends up with stuff being slow.
@@
Jointly considering the available GPU memory with choosing the RAG configuration knob avoids this pitfall. For example, in Figure 8 (b), if the original configuration was stuff, RAGServe can choose to use map_reduce (based on the current GPU memory available).
@@
By doing so, RAGServe can start putting the mappers which fit in memory, into the current batch of requests which fits in the GPU. While map_reduce requires more compute, in this case, it benefits from being able to start execution much faster, as some of the mappers fit in memory.
@@
RAGServe does not need to wait for the GPU memory to free up and changes the configuration jointly with the system resource, to save delay and achieve a better quality-delay tradeoff.
@@
How do we choose the configuration knob’s value jointly?  RAGServe first provides us with a pruned range of configurations. A straw-man solution is to pick a constant value from the across queries. (e.g., the median value of the num_chunks and intermediate_length). While this is better than using one static configuration for all queries, it is still sub-optimal as it does not look at the current system resource availability. This prevents us from exploiting the best quality-delay tradeoff across RAG queries.
@@
We use a best-fit algorithm to allow for variation in configurations across queries. We first compute the GPU memory requirement for the RAG query from the RAG configuration knobs (e.g., num_chunks) for every configuration in the pruned space. Then, we measure the current available memory on the GPU to see what can fit into the current batch.
@@
We then pick the best configuration (with the highest possible knob value) from the pruned space that fits into the GPU. The insight here is higher values of knobs in the pruned space lead to slightly better quality on average. For example, suppose the pruned space says num_chunks is 5-10 and the synthesis_method is stuff, and 5 or 6 chunks can both fit in memory, we choose 6 chunks. We don’t pick a configuration that doesn’t fit, so we would not choose more than 6 chunks. If we do that, the system will queue the query, leading to increased delay.
@@
Once the configuration that fits into the current batch is chosen, the vLLM engine is optimized to perform chunked_prefill. However, even with chunked_prefill, it can only offload parts of long prefill of stuff requests which do not fit in the current batch and still cause additional queuing delay.
@@
What if none of the configurations fit in the GPU?  A main insight for RAGServe’s design comes from the observation that in general, the RAG-specific focused configurations can be loosely-decoupled from the scheduling-specific configurations. RAGServe tries to fit the best possible configurations into GPU memory after it gets the profiler’s reduced configuration space. It can sometimes happen that the current GPU memory availability is too low and none of the profiler’s configurations fit in the current GPU.
@@
One way we handle this is by falling back to a cheaper fixed configuration and choosing to ignore the output of the pruning. As we already have access to the query complexity profile and we can pick cheaper configurations, which would meet the requirement for the current query.
@@
For example, if the query doesn’t require any joint reasoning, we can pick a map_rerank configuration with as many chunks that fit into the current GPU memory, irrespective of what the pruned spaces says.
If joint reasoning is required, we pick a stuff or map_reduce configurations with the few chunks that fit into memory. We can choose which synthesis method to use once based on the exact memory availability.
@@
4.4 RAGServe is just a practical heuristic
It is important to notice that the concept of RAGServe belongs to an emerging trend in the ML and systems community that leverages LLM outputs to guide real system decisions and optimizations, an example of which is LLM routing [53, 30, 13, 50]. While LLM routers use trained LLMs to map decisions from query complexity to choose from families of inference models (outside the realm of RAG), we differ by mapping the output to the configuration knob we run for the RAG queries.
Like these prior efforts, RAGServe is just a heuristic to best utilize the LLM-generated information to guide system optimizations. While it demonstrates remarkable improvement in practice, more work will be needed to complement it for better interpretability and robustness.
5 Refinements to RAGServe
@@
In spite of it all, it is possible for the profiler to (sometimes) fail and in such cases, it is important to detect if RAGServe’s profiler fails on a query in a fast manner to prevent it from leading to bad RAG configurations. Also it is useful to decide how to provide feedback to RAGServe to improve.
When is the quality profile reliable:  RAGServe uses LLM to generate the quality profile. Inspired by recent work in use of model confidence [74, 25, 20] as a quality metric, we use confidence scores for RAGServe’s LLM profiler as to measure the reliability of the profile provided. We obtain the confidence scores from the LLM’s log-probs values on the output (the logarithm of the confidence score, which is directly provided with the output with no extra overhead).
@@
Figure 9: Using a confidence score threshold for different profiler outputs can be used to decide when not to use the output of the profiler.
@@
We then threshold the confidence score using a confidence score threshold ($90\%$ across different datasets) to predict whether the quality profile derived from the quality profiler LLM is actually good (defined as whether the profile can lead to $10\%$ increase in F1-score or 1.52×1.5-2$\times$1.5 - 2 × reduction in delay or both) or not. Such $90\%$ threshold can be tuned for better performance, and we leave it to future work. From Figure 9, we draw two conclusions. First, over 93% of the quality profiles derived from LLM are of high confidence (i.e., over 90%). Further, for those high-confidence profile, over 96% of them are good profiles, meaning that they can be used to improve quality, or reduce latency, or both.
@@
To handle those cases where the quality profile is of confidence score lower than 90%, RAGServe will fall back to the pruned configuration space of recent 10 queries.
@@
How to improve the profiler over time:  RAGServe improve the query profiler LLM by profiling extra feedback prompt to this LLM. We generate this feedback prompt by generating the most accurate output, which is obtained by performing inference on the most resource-demanding configuration (the map_reduce configuration with a large number of input chunks (30) and a high value of intermediate length (300 tokens)). And then ask the quality profiler LLM what configuration it should choose based on the query and the most accurate answer to that query.
@@
The key insight is that, the most accurate answer to the query provides the quality profiler LLM extra knowledge of the knowledge database, and thus can be used to further improve its decision.
@@
To control the cost of generating feedback prompts, RAGServe only generates the feedback prompt once every 30 queries and we only keep the last four feedback prompts.
The cost of RAGServe’s quality profiler:  As for the choice of quality profiler LLM, we mainly use GPT-4o. Though GPT-4o is expensive, as RAGServe only uses it sporadically and only runs it on the query itself, the cost of it is marginal in RAGServe and thus RAGServe still saves cost (as shown in Section 7). We also show that RAGServe can use alternative LLMs as the external LLM to provide feedback and still provides delay reduction without hurting the accuracy in Section 7.
6 Implementation

### v3

@@ 4. METIS: Enabling per-query configuration adaptation for RAG
[…]
@@
    RAG scheduler: Within the pruned configuration space for the query, METIS’ scheduler chooses the best configuration for the query to achieve the best quality-latency trade-off based on the available system resources (§4.3).
Once the configuration is chosen, the METIS’ executes the query using the chosen configuration—retrieving the selected number of chunks and uses the selected synthesis method to feed into the LLM’s input.
Figure 6. METIS consists of a RAG controller which performs configuration space pruning and joint scheduling.
Figure 7. METIS RAG configuration selection workflow.
4.1. Estimating a query’s profile
Query profile: To choose the correct RAG configurations, the first step of METIS is to create the profile of the query (as we see in Figure 7) by querying an LLM (we call this LLM query profiler). We ask the query profiler to estimate four high-level dimensions for each query.
@@ 4.1. Estimating a query’s profile
    Pieces of information required refers to the distinct, standalone pieces of information required to fully answer the query (e.g., the annual income from how many years is required to draw the trend of annual income). The output for this dimension is a number from 1-10.
@@
METIS is not the first to use query profile as a metric for deciding RAG configurations, it extends upon methods like AdaptiveRAG (jeong-etal-2024-adaptive, ) which have used LLM’s to estimate query profile but they only focus on one dimension (the number of chunks to retrieve). However METIS is the first LLM system to introduce resource-quality tradeoff in its RAG decisions with multiple RAG configurations. We show this tradeoff in our experiments, along with the impact of each dimension on the overall improvement, in Section 7.
Why the query profile could be estimated: Estimating the aforementioned query profile is feasible, not only because of the reasoning power of LLMs [^3] We have tested both GPT and Llama models as the profile query-profiler, and they yield similarly impressive results (§7). in analyzing natural language queries, but also because we provide sufficient information to the LLM-based profiler. METIS feeds the profile estimator with not only the query, but also a metadata of the database that contains the background document.
The metadata is a short description about the type of content in the database and its data size (chunk_size). Specifically, we use a single-line summaries already attached to the original source datasets as the metadata of the dataset. For example, the metadata for the KG RAG Finsec’s database (kgrag, ) contains quarterly financial reports and questions of Fortune 500 companies with a chunk_size of 1000. It describes the content topics of the chunks with information such as revenue growth indicators, product release information, sales etc.,. When presented with a query on financials of such a company, the LLM can use the metadata to decide questions like how much to summarize/how much reasoning is required. We give details on the prompt and the intuition to generate metadata for new datasets in Appendix §A.
[…]
@@
Moreover, we observed that extra information does not significantly improve the profiler’s estimates. For instance, in theory, it helps to know the embedding algorithm used by RAG. Yet, the embedding models perform similarly overall across queries and datasets under our consideration. This explains their limited contribution to the profiler, though more future work is needed to understand the wider implications.
@@
[…]
@@
4.2. Mapping query profiles to RAG configurations
After METIS obtains the query profile using the LLM, it performs rule-based mapping to generate values for RAG configuration knobs (e.g., synthesis_method etc. introduced in §2). based on the query profiler’s outputs.
How we map and why the profile helps: To understand the role of query profiles, consider the following examples:
  • $\bullet$
    “Who is the current CEO of NVIDIA?” This query is not complex and does not require joint reasoning. Due to the query being simple with no reasoning required and one piece of information (name of CEO).
  • $\bullet$
    “Which month had the highest NVIDIA’s stock price the six months from January to June 2024?” This query is simple but still needs to read information jointly, specifically six pieces of information (stock price for every month)
  • $\bullet$
    “What are the reasons for NVIDIA’s month-on-month stock price change from January to June 2024” This query is complex and needs to read multiple pieces of information jointly (stock prices, reasons for change etc.) As multiple reasons need to be analyzed here, summarizing all of the information first helps narrow down to relevant information and perform clearer reasoning (why the prices changed).
Algorithm 1 outlines the rule-based mapping process. This mapping is significantly helpful, it improves upon raw profiler outputs and converts them to usable RAG configurations. It reduces the cost of the profiler LLM by restricting it to provide short binary decisions only.
[…]
Algorithm 1 is central to METIS’ design to reduce to the space to our useful RAG configurations and this is extendable to other RAG configurations. For instance, a particular RAG pipeline might use an external re-ranker (Gao_2022, ; ma2023zeroshotlistwisedocumentreranking, ), query re-writer (ma2023queryrewritingretrievalaugmentedlarge, ; jiang2023activeretrievalaugmentedgeneration, ) or perform an external web-search (singh2025agenticretrievalaugmentedgenerationsurvey, ) along with database retrieval. The mapping algorithm can map the profiling LLM’s output (e.g., of Query complexity) and be used to guide such decisions for these newer RAG configurations.
Finally, it is important to note that the concept of METIS belongs to an active research trend in the ML and systems community that leverages LLM outputs and mapping functions to guide real system decisions and optimizations, an example of which is LLM routing (ong2024routellmlearningroutellms, ; hu2024routerbenchbenchmarkmultillmrouting, ; ding2024hybridllmcostefficientqualityaware, ; mohammadshahi2024routoolearningroutelarge, ). While current LLM routers use trained LLMs to map decisions from query complexity to only choose from families of inference models (outside the realm of RAG), we differ by mapping the output to the configuration knob we run for RAG queries.
Like these prior efforts, METIS is a heuristic to best utilize the LLM-generated information to guide system optimizations. While it demonstrates remarkable improvement in practice, more work will be needed to complement it for better interpretability and robustness.
4.3. Joint configuration-scheduling adaptation
Once provided with the narrowed range of each RAG configuration knob (synthesis_method, num_chunks and
@@ 4.3. Joint configuration-scheduling adaptation
intermediate_length), we need to choose a RAG configuration, which is aware of the current system resource (GPU memory). If we pick configurations which do not fit in current memory, it will lead to additional queuing delay waiting for the GPU memory to free up.
@@
We have METIS’s pruned configuration space where the quality is high, we now focus on choosing the best configuration which fits in memory, without focusing on quality.
Figure 8. METIS joint schedules RAG configurations with available GPU memory (chosen example - map_reduce)
@@
Why we need to choose the scheduling jointly: We motivate the need for joint scheduling along with the RAG configuration choice in Figure 8.
@@
Consider a setup where we tune only one RAG configuration knob of synthesis_method. Other knobs num_chunks and intermediate_length are fixed at 20 and 100 respectively. Let’s assume both stuff and map_reduce are present in the pruned space. For the scheduling knob, we consider the amount of GPU memory available for the current batch.
@@
The baseline system in Figure 8 (a) does not consider other jobs in the system and does not evaluate the amount of available resource to make its scheduling decision. Due to its long input length with 20 chunks, stuff turns out to be memory-intensive. If the available GPU memory is low, stuff doesn’t fit in memory and needs to be queued. This ends up with stuff being slow.
@@
Jointly considering the available GPU memory with choosing the RAG configuration knob avoids this pitfall. For example, in Figure 8 (b), if the original configuration was stuff, METIS can choose to use map_reduce (based on the current GPU memory available).
@@
By doing so, METIS can start putting the mappers which fit in memory, into the current running_batch of requests which fits in the GPU. While map_reduce requires more compute, in this case, it benefits from being able to start execution much faster, as some of the mappers fit in memory.
@@
METIS does not need to wait for the GPU memory to free up and changes the configuration aware of system resource, to save delay and achieve a better quality-delay tradeoff.
@@
Jointly choosing the configuration knobs: METIS first provides us with a pruned range of configurations. A straw-man solution is to pick a constant value from the across queries. (e.g., the median value of the num_chunks). While this is better than using one static configuration for all queries, it is still sub-optimal as it does not look at the current system resource availability. This prevents us from exploiting the best quality-delay tradeoff across RAG queries.
@@
We use a best-fit algorithm on underlying vLLM’s continuous batching to allow for variation in configurations across queries. METIS first computes the GPU memory requirement for the RAG query from the RAG configuration knobs for every configuration in the pruned space. For RAG queries, the memory required (e.g., the KVCache size) is measured from the input token length, parameters of the serving model and the quantization (bytes per token). We keep a small 2% buffer size on top of the measurement to deal with potential OOM crashes. We measure the current available memory on the GPU to see what can fit into the current batch.
@@
We then pick the best configuration from the pruned space that fits into the GPU. METIS defines the best configuration as the one with overall highest memory requirement, from all which fit in memory. The insight here is that within the reduced range of good quality configurations, higher memory configurations correspond to expensive configurations (e.g. more number of chunks, higher intermediate length). In general, these configurations should lead to slightly higher quality in the reduced space. For example, if the pruned space says num_chunks is 5-10 and the synthesis_method is stuff and both 5 or 6 chunks can fit in memory, we choose 6 chunks. We don’t pick a configuration that doesn’t fit in GPU, so we would never choose more than 6 chunks. If we do that, the system will queue the request inflating the delay.
@@
After choosing the configuration that fits into the current running_batch, the vLLM engine is optimized to perform chunked_prefill. However, even with chunked_prefill, it can only offload parts of long prefill of stuff requests which do not fit in the current batch and still inflates the queuing delay. Jointly scheduling RAG configurations enables efficient resource usage, which cannot be obtained by only relying on the output of the LLM profiler.
@@
What if none of the configurations fit in the GPU? A main insight for METIS’s design comes from the observation that in general, the RAG-specific focused configurations can be loosely-decoupled from the scheduling-specific configurations. METIS tries to fit the best possible configurations into GPU memory after it gets the profiler’s reduced configuration space. It can sometimes happen that the current GPU memory availability is too low and none of the profiler’s configurations fit in the currently available GPU.
@@
METIS handles this is by falling back to a cheaper fixed configuration and ignoring the output space of the pruned configurations. As METIS already have access to the query complexity profile, we can pick from cheaper configurations, which would meet the requirement for the current query.
@@
If the query does not require joint reasoning, we pick a map_rerank with as many chunks that fit into available GPU memory. If joint reasoning is required, we pick a stuff with as many chunks that fit into memory. METIS does not queue a configuration from outside the pruned range if none fit, but falls back to a fitting configuration just outside the range.
@@
5. Refinements to METIS
@@
In spite of it all, it is possible for the profiler to (sometimes) fail and in such cases, it is important to detect if METIS’s profiler fails on a query in a fast manner to prevent it from leading to bad RAG configurations. Also it is useful to decide how to provide feedback to METIS to improve.
@@
Figure 9. Confidence score threshold for different profiler outputs is used to decide when not to use the profiler output.
When is the quality profile reliable? METIS uses LLM to generate the quality profile. Inspired by recent work in use of model confidence (xiong2024can, ; geng-etal-2024-survey, ; 10.1145/3637528.3671470, ) as a quality metric, we use confidence scores for METIS’s LLM profiler as to measure the reliability of the profile provided. We obtain the confidence scores from the LLM’s log-probs values on the output (the logarithm of the confidence score, which is directly provided with the output with no extra overhead).
@@
We then threshold the confidence score using a confidence score threshold ($90\%$ across different datasets) to predict whether the quality profile derived from the quality profiler LLM is actually good (defined as whether the profile can lead to $10\%$ increase in F1-score or $1.5-2\times$ reduction in delay or both) or not. Such $90\%$ threshold can be tuned for better performance, and we leave it to future work. From Figure 9, we draw two conclusions. First, over 93% of the quality profiles derived from LLM are of high confidence (i.e., over 90%). Further, for those high-confidence profile, over 96% of them are good profiles, meaning that they can be used to improve quality, or reduce latency, or both.
@@
To handle those cases where the quality profile is of confidence score lower than 90% , METIS will fall back to the pruned configuration space of recent 10 queries.
@@
How to improve the profiler over time? METIS improves the query profiler LLM by profiling extra feedback prompt to this LLM. We generate this feedback prompt by generating the most accurate output, which is obtained by performing inference on the most resource-demanding configuration (the map_reduce configuration with a large number of input chunks (30) and a high value of intermediate length (300 tokens)) and then ask the quality profiler LLM what configuration it should choose based on the query and the most accurate answer to that query.
@@
The key insight is that, the most accurate answer to the query provides the quality profiler LLM extra knowledge and thus can be used to further improve its decision.
@@
To control the cost of generating feedback prompts, METIS only generates the feedback prompt once every 30 queries and we only keep the last four feedback prompts.
[…]
6. Implementation